### v1

# RV-Syn: Rational and Verifiable Mathematical Reasoning Data Synthesis based on Structured Function Library

Abstract
The advancement of reasoning capabilities in Large Language Models (LLMs) requires substantial amounts of high-quality reasoning data, particularly in mathematics. Existing data synthesis methods, such as data augmentation from annotated training sets or direct question generation based on relevant knowledge points and documents, have expanded datasets but face challenges in mastering the inner logic of the problem during generation and ensuring the verifiability of the solutions. To address these issues, we propose RV-Syn, a novel Rational and Verifiable mathematical Synthesis approach. RV-Syn constructs a structured mathematical operation function library based on initial seed problems and generates computational graphs as solutions by combining Python-formatted functions from this library. These graphs are then back-translated into complex problems. Based on the constructed computation graph, we achieve solution-guided logic-aware problem generation. Furthermore, the executability of the computational graph ensures the verifiability of the solving process. Experimental results show that RV-Syn surpasses existing synthesis methods, including those involving human-generated problems, achieving greater efficient data scaling. This approach provides a scalable framework for generating high-quality reasoning datasets.
RV-Syn: Rational and Verifiable Mathematical Reasoning
Data Synthesis based on Structured Function Library
Jiapeng Wang1thanks:   Equal contribution., Jinhao Jiang [^1] footnotemark: 1, Zhiqiang Zhang, Jun Zhou, Wayne Xin Zhao1thanks:   Corresponding author., 1Gaoling School of Artificial Intelligence, Renmin University of China 2Ant Group wangjp1010@ruc.edu.cn, batmanfly@gmail.com
1 Introduction
The development of advanced reasoning Large Language Models (LLMs) Zhao et al. (2023); OpenAI (2024) has markedly improved their ability to address complex tasks across domains such as mathematics, science, and coding. This highlights the importance of synthesizing complex reasoning data to drive further advancements, given the limited availability of high-quality annotated instructions Shah et al. (2024); Yuan et al. (2023).
To address this scarcity, researchers have explored various synthesizing methods, particularly in the mathematics domain. The mainstream methods involve data augmentation based on existing annotated training sets, such as GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021), ranging from self-evolving instructions Xu et al. (2024); Zeng et al. (2024) and question paraphrasing Yu et al. (2024), to solution augmentation Lu et al. (2024). However, these methods are limited by the available training data, constraining the synthesis diversity Li et al. (2024c, 2023). To enhance diversity, recent approaches enable LLMs to generate a large scale of questions from various mathematics-related sources, including web pages Yue et al. (2024) and knowledge points Tang et al. (2024) from web corpora or textbooks.
Figure 1: (Left) Illustration of the limitations of existing synthesis methods; (Right) Examples of data synthesized by RV-Syn, which naturally support sophisticated reasoning structures and provide ground-truth answers.
However, as shown in Figure 1, these direct problem generation methods encounter challenges in lacking a deep understanding and simulation of the underlying reasoning processes. As a result, the synthesized problems may appear superficially plausible but actually suffer from insufficient reasoning depth and logical inconsistencies (such as internal contradictions), especially when generating complex multi-step reasoning problems. Furthermore, a long-standing challenge in synthetic data generation is the difficulty in supporting the validation of solution correctness during the synthesis process, ultimately compromising training efficiency.
To address the aforementioned challenges, we draw inspiration from how human educators create problems. Just as one cannot directly write down an Olympic-level math problem without deep consideration and careful curation of the underlying problem-solving process, the same holds true for models. Specifically, human educators obtain abstract independent computation goals from historical experience, such as “projecting a vector onto a plane”, and “solving for the smallest prime number that satisfies specific conditions”. Then, when designing a new question, they would combine these computation goals to obtain a new solution or introduce them into an existing solution. Finally, they derive the final problem based on this elaborated solution process. This strategy ensures the development of problems with coherent and logically consistent solution processes, thereby minimizing the internal inconsistencies of the final problems.
Inspired by this, we propose RV-Syn, a novel method for mathematical problem synthesis that emphasizes enhanced rationale and verifiability. This method comprises three stages. First, we extract solution processes from existing problems by prompting the model to generate code-based solutions (computational graphs) where each function performs a specific mathematical operation. Second, a structured function library in graph format is built, detailing the relationships between these functions. Third, a graph-aware sampling strategy combines these functions to create new computational graphs, which are then executed and back-translated to synthesize new problems with correct answers. By generating questions through solution guidance and utilizing the structured and executable nature of code as a blueprint for problem design, RV-Syn effectively provides the rich semantics and control flows required for complex reasoning. As shown in Figure 1, this enables the synthesis of diverse and sophisticated reasoning structures—often challenging to achieve for direct problem generation—including multi-hop reasoning (sequential execution), iterative loops (for-loops), and modular structures (function calls), while naturally providing ground-truth labels.
We conducted extensive experiments to evaluate the proposed RV-Syn method using various LLMs. The results demonstrate that our method achieves superior average performance across five benchmarks compared to existing approaches, including those involving human-designed problems (e.g., Numina-Math). Notably, RV-Syn outperforms the previous state-of-the-art method while utilizing only half the data, leading to a more efficient scaling curve.
2 Related Work
Mathematical Reasoning.
Researchers have proposed various approaches to enhance the mathematical reasoning capabilities, including methods applied in training or inference stages. During the training stage, existing methods aim to enhance the LLMs from the aspects of pre-training and post-training. Specifically, some studies Shao et al. (2024); Chen et al. (2024a) involve collecting extensive math-related corpora and enhancing the foundational mathematical capabilities of LLMs through continual pre-training. In contrast, other studies Tong et al. (2024); Huang et al. (2024b); Toshniwal et al. (2024) focus on synthesizing a substantial amount of high-quality math-related instructional data, further refining the problem-solving abilities of LLMs through post-training. During the inference stage, there are two primary prompting approaches, including Chain-of-Thought (CoT) Wei et al. (2022) and Program-of-Thought (PoT) Chen et al. (2023). Specifically, some studies Yao et al. (2023); Shinn et al. (2023) utilize CoT to elicit LLMs’ inherent reasoning ability with a step-by-step reasoning process. Furthermore, other studies Gou et al. (2024); Wang et al. (2024); Chen et al. (2024b); Wen et al. (2024) incorporate PoT to allow LLMs to utilize external computors during reasoning. In this work, we focus on the training stage and try to synthesize instruction data by leveraging the strengths of both the CoT and PoT methods.
Data Synthesis for Math Instruction.
Existing research has proposed various synthesizing approaches: Firstly, existing studies focus on synthesizing new mathematical problems based on few-shot problems Li et al. (2024a), math-related documents Yue et al. (2024), key knowledge points Huang et al. (2024a), or self-evolution Luo et al. (2023). Furthermore, with the improvement of LLMs’ capabilities, subsequent studies Zhou et al. (2024); Ding et al. (2024) explore to utilize of open-source LLMs (i.e., DeepSeek-Math and Qwen2-Math) to scale the number of synthesized problems. Despite their effectiveness, direct problem generation suffers from a lack of fine-grained control over the problem’s internal logic, leading to lower-quality problems. Besides, these methods frequently lack mechanisms for validating the correctness of generated answers, which may introduce errors and affect the final performance after instruction tuning. In contrast, our method starts from the internal logic of problem-solving, rather than directly synthesizing problems. This provides more refined logical consistency and improves problem quality, while also enabling effective correctness control.
Figure 2: The pipeline of our proposed method. The upper part demonstrates the construction of the operation function library. The lower part illustrates the synthesis process.
3 Approach
In this section, we provide a comprehensive introduction to our proposed RV-Syn method. Its essence involves developing a comprehensive set of mathematical operations and skills. These skills are then systematically combined to construct the solution process, ultimately enabling the generation of complex problems through back-translation. The complete pipeline comprises three core stages: (1) decomposing existing seed problems to generate a set of computational graphs (Section 3.1), (2) extracting functions from these computational graphs and constructing a graph-format function library (Section 3.2), and (3) combining selected functions to reconstruct new computational graphs, which are subsequently back-translated into problems (Section 3.3). The overall data flow of our method is illustrated in Figure 2.
3.1 Problem Decomposition
We first represent the solutions of existing mathematical problems with sequences of functions into a standardized format shown in Appendix E leveraging LLMs. Each function represents a specific mathematical skill with defined parameters, forming an executable computational graph that captures the data flow of the problem-solving process (examples shown in Figure 2). To ensure the correctness of the extracted solutions and their associated functions, we filter out erroneous cases by executing the solution code using a Python interpreter and comparing its output with the ground-truth answer provided in the corresponding Chain-of-Thought annotation. This results in a large-scale, high-quality set of computational graphs derived from real-world problems.
3.2 Graph-based Function Library Construction
After decomposing the problem, we extract functions from the computational graphs. To further enhance the quality of the function library, we employ model-based labeling. We utilize LLMs to validate the correctness of the functions and identify their corresponding mathematical topics, eliminating any functions that contain errors. The topic definitions are provided in Table 6, and the prompts are shown in Appendix C. To describe the relationships between these functions, each function is represented as a node, and they are organized into a graph-based function library.
Specifically, the construction process involves two key steps: structure and semantic-aware function merging, followed by co-occurrence and topic-based node connecting.
3.2.1 Structure and Semantic-aware Function Merging
As previously discussed, each extracted function primarily encapsulates a specific mathematical operation or skill. Consequently, different functions may exhibit similarities in semantics (i.e., possessing similar skills) or structure (i.e., sharing similar computational logic). To preserve the diversity of function expressions while minimizing redundancy within the final function library, we merge functions that are similar in either structure or semantic aspects into one node. Therefore, the final node is a set that contains one or more similar nodes. In the subsequent description, we will take this setting as the default.
Specifically, we examine the similarity between two functions by analyzing both structure and semantic aspects. For the structure, we focus on the computational operations without considering specific variable names. For instance, expressions such as “$(a+b)/c$” and “$(m+n)/d$” are considered equivalent. To accomplish this, we extract the function body and parse it using Abstract Syntax Trees (ASTs). For the semantic, we compute hash values for the function’s docstrings. If the ASTs or hash values of two functions are identical, we merge these functions into a single node. We show an example of two similar functions in Table 9 at Appendix D. This approach allows similar functions (e.g., matrix multiplication and matrix projection) to preserve their distinct mathematical meanings while minimizing redundancy in the final function library, offering advantages over direct deduplication.
3.2.2 Co-occurrence and Topic-based Node Connecting
After obtaining the final set of nodes through function merging, we further establish connections between relevant nodes to construct the final graph. We define two types of edges: co-occurrence-based edges and topic-consistent edges.
Specifically, for any two given nodes (i.e.,, two sets of functions), we first check whether there exists at least one pair of functions—one from each set—that co-occurred in the same computational graph. If so, we add a co-occurrence-based edge between these two nodes. If not, we then check whether there is at least one pair of functions sharing the same topic. If this condition is met, we add a topic-consistent edge. If neither condition is satisfied, no edge is added between the nodes.
In this way, we finally obtain an incompletely connected graph that contains multiple types of edges, which depicts the relationships between nodes. This serves as the basis for the subsequent recombination of functions. Some statistics of the graph-based function library are shown in Table 6 and 7 in Appendix A.
3.3 Problem Synthesis
Based on the constructed graph-format function library, we are able to synthesize complex problems in a solution-guided logic-aware approach. This process consists of two primary steps: first sampling various functions to regenerate the computational graph, then synthesizing the final complex problems.
3.3.1 Computational Graph Regeneration
To balance the reasonableness and novelty of the computational graphs, we design three sampling strategies: first, selecting nodes connected by co-occurrence-based edges; second, selecting nodes connected by topic-based edges; and third, selecting unconnected nodes. During the sampling process, nodes are randomly selected using these strategies. Since each node represents a set of functions, we further randomly select one function per node. Subsequently, we generate a computational graph based on these selected functions as detailed in Appendix C.
This approach offers several advantages: Firstly, by combining functions, the generated computational graph can comprehensively cover mathematical skills while minimizing redundancy. Secondly, the sampling strategy effectively balances high-frequency combinations of mathematical skills (i.e., connected by edges) with long-tail combinations (i.e., without connected edges). In our setup, we randomly select from the three strategies when synthesizing each data sample.
3.3.2 Problem Back-translation
After obtaining the new computational graphs, we execute all computational graphs and filter out those that cannot be executed effectively, thereby obtaining answers for the valid computational graphs. Subsequently, we utilize LLMs to generate problems based on these graphs, as detailed in Appendix C. This method is similar to the approach human educators use when creating questions: they first consider the solution process and then formulate corresponding questions. Consequently, this approach enhances the grounding and logical coherence of the final synthesized problems.
Method Category Open-source Problem Crafter Rational Problem Synthesize Automatic Solution Verification
MetaMath Aug
Orca-Math Aug
WizardMath Aug
MathGenie Aug
Mammoth2 New
Jiuzhang3.0 New
ControlMath New
MathScale New
PromptCoT New
ScaleQuest New
RV-Syn New
Table 1: Comparison of Different Methods. Category specifies whether the method augments existing data (“Aug”) or synthesizes new questions (“New”). Open-source Problem Crafter indicates whether the method utilizes open-source models (✓) or proprietary models like GPT-4 (✗) for problem generation. Rational Problem Synthesize denotes whether the synthesized problems incorporate internal problem-solving logic. Automatic Solution Verification denotes the method’s ability to automatically verify the correctness of synthesized data.
3.4 Comparison to Previous Work
We give a comparison in Table 1. The first line of research, including MetaMath Yu et al. (2024), Orca-Math Mitra et al. (2024), WizardMath Luo et al. (2023), MuggleMath Li et al. (2024b), MathGenie Lu et al. (2024), primarily focuses on augmenting existing problems or solutions, which leads to new instructions that are too similar to the original ones, thus limiting diversity. Another line of research, represented by Mammoth2 Yue et al. (2024), Jiuzhang3.0 Zhou et al. (2024), KPMath Huang et al. (2024a), and ScaleQuest Ding et al. (2024), synthesizes brand new mathematical problems based on math-related documents, knowledge points, or from scratch, offering greater diversity and scalability. However, many of these methods overlook the internal logical processes when directly synthesizing problems. Methods like ControlMath Chen et al. (2024c) and PromptCoT Zhao et al. (2025) attempt to enhance the rationality of synthesis by leveraging equation generators or injecting intermediate thought processes. Nevertheless, they still lack mechanisms to verify the correctness of the synthesized data. Our RV-Syn method approaches problem generation from the perspective of internal problem-solving logic rather than directly producing problem statements without methodical consideration, thereby enhancing problem quality. Additionally, through executable computation graphs, our method enables automatic verification of solution correctness.
Method Problem Crafter MATH-500 GSM8K GSM-Hard College Math Olympiad Bench Avg (Rel. Imp.)
Models based on LLaMA-3-8B-Instruct
Official Model - 28.4 75.1 35.6 21.2 7.3 33.50
MetaMath ChatGPT 42.4 83.1 36.2 22.3 10.1 38.82 (+15.9%)
Orca-Math GPT-4 29.4 86.6 42.7 19.8 11.1 37.92 (+13.2%)
Mammoth2 72B Model 42.2 78.6 40.3 33.0 13.6 41.54 (+24.0%)
Jiuzhang3.0 7B Math Model 43.4 81.3 40.5 29.6 15.9 42.14 (+25.8%)
MathScale ChatGPT 43.4 81.3 38.4 32.6 13.6 41.86 (+25.0%)
PromptCoT 72B Model 43.4 78.2 41.5 25.8 15.7 40.92 (+22.1%)
ScaleQuest 7B Math Model 45.8 83.6 41.9 29.4 13.3 42.80 (+27.8%)
MMIQC Human & GPT-4 46.6 83.5 41.2 30.6 16.4 43.66 (+30.3%)
Numina-Math Human & GPT-4 49.8 81.8 40.3 31.5 16.1 43.90 (+31.0%)
RV-Syn 7B / 72B model 50.4 82.6 44.5 30.7 16.4 44.92 (+34.1%)
Models based on Qwen2.5-7B-Instruct
Official Model - 74.0 86.8 59.9 44.1 36.0 60.16
MetaMath ChatGPT 74.0 90.9 67.6 45.5 35.0 62.60 (+4.1%)
Orca-Math GPT-4 74.2 92.0 67.4 45.3 35.4 62.86 (+4.5%)
Mammoth2 72B Model 75.0 90.2 68.2 45.9 34.7 62.80 (+4.4%)
Jiuzhang3.0 7B Math Model 74.8 91.3 67.4 45.8 34.7 62.80 (+4.4%)
MathScale ChatGPT 74.0 91.0 67.6 45.3 36.4 62.98 (+4.5%)
PromptCoT 72B Model 75.8 90.5 67.7 45.8 34.8 62.92 (+4.6%)
ScaleQuest 7B Math Model 75.0 91.4 67.6 45.9 34.4 62.86 (+4.5%)
MMIQC Human & GPT-4 75.8 91.1 68.0 45.6 37.0 63.50 (+5.5%)
Numina-Math Human & GPT-4 75.4 91.7 68.5 46.3 36.9 63.76 (+5.9%)
RV-Syn 7B / 72B model 76.8 91.3 69.6 46.1 36.4 64.04 (+6.4%)
Models based on Phi-3-mini
Official Model - 43.0 73.3 54.1 35.7 15.7 44.40
MetaMath ChatGPT 55.8 89.8 64.5 39.6 20.7 54.08 (+21.8%)
Orca-Math GPT-4 55.2 90.4 65.3 40.0 19.6 54.10 (+21.9%)
Mammoth2 72B Model 59.0 87.7 62.9 41.6 21.3 54.50 (+22.7%)
Jiuzhang3.0 7B Math Model 58.2 88.9 63.8 41.5 19.6 54.40 (+21.8%)
MathScale ChatGPT 58.2 89.5 63.5 40.3 20.3 54.36 (+22.5%)
PromptCoT 72B Model 58.2 87.4 63.5 39.6 21.3 54.00 (+21.6%)
ScaleQuest 7B Math Model 59.0 88.9 63.8 41.2 21.8 54.94 (+23.7%)
MMIQC Human & GPT-4 59.4 89.1 64.3 41.8 21.8 55.28 (+24.5%)
Numina-Math Human & GPT-4 60.0 88.8 63.8 42.2 21.9 55.34 (+24.6%)
RV-Syn 7B / 72B model 59.8 88.8 65.0 42.2 22.4 55.64 (+25.3%)
Table 2: After training on 50k data, the evaluation results of our method compared with various synthesis approaches and two datasets containing human-crafted problems across five benchmarks. The best and second-best ones among LLMs with the same backbone model are marked in bold and underlined respectively.
4 Experiments
4.1 Experimental Setup
Compared Baselines.
We compare our method with previous direct problem synthesis methods with publicly available datasets, including: (1) MetaMath Yu et al. (2024); (2) Orca-Math Mitra et al. (2024); (3) MathScale Tang et al. (2024);(4) Mammoth2 Yue et al. (2024); (5) Jiuzhang3.0 Zhou et al. (2024);(6) PromptCoT Zhao et al. (2025) and (7) ScaleQuest Ding et al. (2024). We also add two datasets containing human-crafted problems as strong baseline: (7) MMIQC Liu et al. (2024) contains QA pairs from Mathematics Stack Exchange and iteratively constructs more complex problems from MATH training data; (8) Numina-Math LI et al. (2024) is a large collection of both real-world human-crafted and synthesized math problems. We utilize the high-quality cn_k12, Olympiads, and synthetic_math subsets.
Tuning Data.
Directly comparing the performance of models from previous works introduces potential unfair factors, as newer models may leverage more advanced models for backbone and answer generation. To ensure a fair comparison, we use the same model, Qwen-2.5-Math-7B-Instruct Yang et al. (2024) with temperature set to 0.7 and top-$p$ set to 0.95, to generate answers for problems synthesized by each method and train the same backbone model. By controlling for other variables, we can focus solely on comparing the quality of the synthesized problems.
Evaluation and Metrics.
We assess the models’ performance on MATH-500 Hendrycks et al. (2021); Lightman et al. (2024), GSM8K Cobbe et al. (2021), GSM-Hard Gao et al. (2023), College Math Tang et al. (2024) and OlympiadBench He et al. (2024). The generated outputs are all in the form of natural language Chain-of-Thought (CoT) Wei et al. (2022) reasoning through greedy decoding, and we report zero-shot pass@1 accuracy.
Implementation Details.
We extract seed data from the high-quality Numina-Math dataset LI et al. (2024) to collect function nodes and fine-tuning data. The problem decomposition is performed using Qwen2.5-72B-Instruct, resulting in a function library containing 50k nodes. We then use the fine-tuned Qwen2.5-7B-Instruct to generate computational graphs with the strategy described in 3.3.1, followed by problem back-translation using Qwen2.5-72B-Instruct.
4.2 Main Results
As shown in Table 2, methods focusing on synthesizing diverse new problems (e.g., Mammoth2, Jiuzhang3.0, ScaleQuest) outperform approaches that augment existing data (e.g., MetaMath, Orca-Math). This performance gap may be attributed to augmentation-based methods leading to new instructions that are too similar to the original ones, emphasizing the crucial role of data diversity. On the other hand, datasets containing real-world human-designed problems exhibit a clear performance advantage over existing synthetic data, highlighting the superior quality of human-designed problems. Our approach outperforms traditional data synthesis methods and yields results that closely match or even surpass the effectiveness of human-designed problems, without relying on proprietary models like GPT-4. Notably, this advantage is particularly evident on more challenging or computation-intensive datasets such as MATH-500, GSM-Hard, and OlympiadBench, further emphasizing the benefits of rational synthesis.
4.3 Further Analysis
Figure 3: The performance of the model as the training data size increases. The "Average" refers to the mean performance across five datasets: MATH-500, GSM8K, GSM-Hard, College Math, and OlympiadBench.
Scalability Study.
In this section, we study the scalability of our method while maintaining high-quality problem generation. We compare our method with previous approaches by training on scaling datasets using LLaMA-3-8B-Instruct AI@Meta (2024), and further scale our method to 500k to explore its scalability at larger scales. The results in Figure 3 demonstrate a strong scaling curve, with performance consistently improving as the dataset size increases. Our method consistently outperforms other methods, and notably, achieves better performance with only half the dataset size (50k vs. 100k). This advantage is particularly pronounced on challenging or computation-intensive datasets, demonstrating the superiority of our method for enhancing advanced mathematical reasoning capabilities.
Answer Annotation Model Matters.
Previous works often directly fine-tune models on publicly available datasets released by baseline methods and compare their performance. However, we find that the choice of the answer annotation model is one of the most critical factors influencing downstream performance, which introduce substantial unfairness into such comparisons. To illustrate this, we re-annotate existing datasets using a unified model,Qwen2.5-Math-7B-Instruct,and compare the performance of LLaMA-3-8B-Instruct fine-tuned on 50k these newly annotated datasets versus the original versions. The results in Table 3 show that many methods achieve significant performance improvements, and the performance gap between them narrows. In all our experiments, we use the same annotation model to ensure a fair comparison and to isolate the impact of problem quality. We observe consistent performance improvements from our method. Although some of the improvements may appear modest under the uniform annotation setup, they in fact represent substantial enhancements in data quality, which are partially obscured by the controlled evaluation conditions.
Dataset Original Re-annotation Performance Change
Jiuzhang3.0 30.5 42.1 +38.0%
MathScale 26.5 41.9 +58.1%
MMIQC 34.9 43.7 +25.2%
NuminaMath 38.7 43.9 +13.4%
Table 3: Performance of LLaMA-3-8B-Instruct Fine-tuned on Re-annotated Datasets using Qwen2.5-Math-7B-Instruct.
Analysis of Problem Quality.
In this part, we quantitatively analyze the difficulty and efficiency of the synthesis problem using three key metrics, following PromptCoT Zhao et al. (2025). First, accuracy serves as an indicator of problem difficulty. Lower performance signifies a higher level of difficulty. For datasets without ground-truth labels, we adopt the approach detailed in PromptCoT Zhao et al. (2025) by leveraging a more powerful model, Qwen2.5-Math-72B-Instruct, to generate reference answers, applying self-consistency with 8 rollouts to enhance reliability. Second, the average number of reasoning tokens provides insights into the complexity of the reasoning process. A larger average number of tokens suggests that the model necessitates more extensive reasoning steps. Finally, the average accuracy on the benchmark after tuning reflects the dataset’s contribution to improving model performance. Larger improvements in benchmark accuracy observed after fine-tuning suggest that the respective dataset is more effective in enhancing the model’s capabilities. Experimental results in Table 4 show a strong consistency among these three metrics, and our method achieves the best results across all three difficulty evaluation metrics.
Dataset Accuracy ($\downarrow$) Avg. Reasoning Tokens ($\uparrow$) Avg. Benchmark Accuracy ($\uparrow$)
MetaMath 97.2 1,339 38.82 (+15.9%)
Orca-Math 86.3 919 37.92 (+13.2%)
MathScale 82.0 2,565 41.86 (+25.0%)
Mammoth2 70.6 3,272 41.54 (+24.0%)
ScaleQuest 75.8 2,591 42.80 (+27.8%)
NuminaMath 72.9 3,858 43.90 (+31.0%)
RV-Syn 55.6 4,431 44.92 (+34.1%)
Table 4: Difficulty and efficacy evaluation for different mathematical datasets. Accuracy: Performance of Qwen2.5-Math-7B-Instruct on the problems in different datasets. Avg. Reasoning Tokens: Average number of tokens in reasoning processes generated by DeepSeek-R1-Distill-Qwen-7B when processing the problems in different datasets. Avg. Benchmark Accuracy: Performance of LLaMA-3-8B-Instruct after fine-tuning on different datasets.
Correctness Verification.
In this part, we investigate the verifiable characteristics of our method to ensure data correctness. We sample 1,000 data points from the training data of each method and evaluate them using GPT-4o. We validate this evaluation method in Appendix B and analyze the error rates for both problems and solutions.
Method Problem Error Rate (%) Solution Error Rate (%)
Mammoth2 0.9 10.6
Jiuzhang3.0 2.4 6.5
MathScale 1.6 4.9
ScaleQuest 5.6 8.0
MMIQC 1.3 7.6
Numina-Math 0.6 5.5
RV-Syn 0.9 1.4
Table 5: Error rate analysis of training data across different methods. The problem error refers to issues such as inconsistency or conflict in problem conditions and statements; the solution error indicates computational mistakes or misinterpretation of the problem in the solution.
As shown in Table 5, firstly, the problem error rate reflects the quality of the synthesized problems. It can be observed that problems generated by humans exhibit a lower error rate, while directly synthesized problems contain more errors. Our method approaches the level of control that humans exert in problem design. Secondly, the solution error rate reflects the extent to which each method controls the correctness of training data. Since other methods cannot naturally obtain the ground truth, their solution error rates are relatively high. Further quality control requires heavy reliance on powerful models for post-filtering, annotation, voting, or scoring, which incurs significant additional costs. In contrast, our method inherently provides the ground truth and enables rule-based matching, ensuring an extremely low error rate. We then manually inspect our data labeled as erroneous and find that most solutions have correct final answers, with only minor errors in intermediate steps.
Figure 4: Performance comparison of distilling Long-CoT on 20k data.
Distilling Long-CoT.
In addition to short-CoT, we also explore the ability of our data to distill Long-CoT abilities. We replace the answer generation model with DeepSeek-R1-Distill-Qwen-7B Guo et al. (2025) and train Qwen2.5-7B-Instruct on a 20k dataset, comparing our method with ScaleQuest. During the evaluation, we set the maximum sequence length to 16k and exclude samples where both models exceed the maximal length. We show the results in Figure 4. We can see that our synthesis method not only performs well on short-CoT but also excels in distilling Long-CoT compared to the baseline.
5 Conclusion
In this paper, we introduced RV-Syn, a novel approach to mathematical problem synthesis that enhances both rationale and verifiability. It first constructed a graph-format function library, which was composed of a large amount of mathematical functions and was scalable to insert new ones. Then, it generated the new solutions by combining functions into computation graphs in a controllable manner, which can be then executed to support solution verification. Finally, it back-translated these computation graphs to problems. Our experiments demonstrate that RV-Syn outperforms existing methods, achieving higher accuracy with less data. This efficiency highlights its potential for improving the reasoning capabilities of LLMs while minimizing resource consumption. RV-Syn represents a significant advancement in reasoning data generation, offering a scalable and reliable framework. Future work can explore extending it to other reasoning domains and further refining control mechanisms for broader applicability.
Limitations
Despite the promising results demonstrated by the RV-Syn method, several limitations remain. First, the approach relies heavily on the quality and diversity of the initial function library. While we have curated a comprehensive set of functions from a large scale of existing synthetic datasets, there may still be gaps in representing certain mathematical concepts, potentially limiting the scope of problems that can be synthesized. Second, although our experiments show improved performance across several benchmarks, the generalization of RV-Syn to other domains beyond mathematics remains unexplored, such as the code domain. Future work should investigate the adaptability of this approach to different types of reasoning tasks and datasets. We aim to address these limitations in future research by expanding the function library, optimizing computational efficiency, and exploring cross-domain applications.
Appendix A Function Library Statistic
Here, we show the function library statistic in Table 6 and Table 7.
Topic Proportion
Number Theory 0.204
Prealgebra 0.082
Intermediate Algebra 0.080
Algebra 0.102
Precalculus 0.290
Counting & Probability 0.119
Geometry 0.121
Miscellaneous 0.003
Table 6: Topic proportions in function library. Topics for each function were labeled according to the categories in the MATH dataset.
Metric Value
CB Connected Components Number 23419
CB Largest Connected Component Size 10847
CB Average Degree per Node 1.225
TB Connected Components Number 8
TB Largest Connected Component Size 13638
TB Average Degree per Node 8607
Table 7: Graph-based function library statistics. CB denotes Co-occurrence Based, TB denotes Topic Based.
Appendix B Correctness Evaluation Validation
Considering the substantial cost associated with human evaluations, we use one of the most powerful models currently available, GPT-4o, to conduct correctness evaluation. To ensure its validity, we conducted an experiment using GPT-4o to evaluate the accuracy of Qwen2.5-Math-7B-Instruct on the benchmark test set and compared the results with ground-truth evaluations. As shown in Table 8, the evaluation results—74.25% (GPT-4o) vs. 76.61% (Ground truth)—demonstrate a close alignment. False positive samples account for only 1.89% of all cases, while the false negative rate is even higher, indicating that GPT-4o tends to be more stringent in its evaluation. This is likely because it not only verifies the final answer but also checks intermediate steps. Considering that even ground-truth-based evaluations may contain some degree of error, we believe this discrepancy falls within an acceptable range. The prompt we use is shown in Appendix C.
Ground Truth
Correct (+) Incorrect (-) Total
GPT-4o Correct (+) 0.7236 0.0189 0.7425
Incorrect (-) 0.0425 0.2150 0.2575
Total 0.7661 0.2339 1.0000
Table 8: Confusion matrix for GPT-4o evaluation.
Appendix C Prompts
In this part, we present the prompts used in the three main components of our method: Problem Decomposition, Computational Graph Regeneration, and Problem Back-translation.
Prompt for Problem Decomposition Your task is to solve the following math problem using Python code, employing a ‘main‘ function along with several custom operation functions. Problem: {PROBLEM} Answer the question step by step first, understand the logic of the problem. Then: 1. Represent calculation operations as functions, where each function specifies its functionality, input parameter types, and output return types. 2. Each operation function should represent a complete, meaningful computational skill or operation, and should be reusable across different problems. 3. Simple operations such as addition, subtraction, multiplication and division do not need to be expressed as functions. 4. Each function should have a specific, independent purpose for performing a particular calculation. You may use functions from standard or imported libraries (like SymPy, math, etc.; Import within the function itself), but **cannot** use other custom operation functions. 5. The operation functions should be flexible, not hardcoded, and able to handle various scenarios by accepting different parameters. 6. Solve the problem by calling the operation functions within the main function. 7. Only print the answer without any text.
Prompt for Function Labelling Please analyze the following mathematical function operation, judge its functional correctness and assign it a subject category. {function} - First analyze the function and evaluate its correctness. Out put Yes or No for correctness judgement. - There are 7 possible subject categories to choose from: Prealgebra, Intermediate Algebra, Algebra, Number Theory, Precalculus, Geometry, Counting & Probability, and Miscellaneous. Output example: <analyse>analyze here</analyse> <correctness>Yes</correctness> <subject>Number Theory</subject>
Prompt for Computational Graph Regeneration Fine-tuning Given a set of mathematical operation functions: ``` {FUNCTIONS} ``` Write a ‘main‘ function that intelligently combines these operations to solve a meaningful and non-trivial mathematical problem.
Prompt for Problem Back-translation Given a piece of code that solves a mathematical problem: ``` {CODE} ``` First, understand what the code snippet is doing within the ‘<analysis></analysis>‘ tags. Then, identify the original problem the functions are solving and output it within the ‘<problem></problem>‘ tags. Finally, rewrite the problem concisely, in a manner similar to questions found in textbooks or test papers, and output it within the ‘<final problem></final problem>‘ tags. The question should have only one answer.
Prompt for Correctness Verification Your goal is to determine if the given problem and its solution contain mistakes. If both the problem and the solution are correct, output ’Correct’. If the problem contains error such as inconsistency, conflict or no solution, output ’Problem Error’. If the problem is correct but the solution contains error, output ’Solution Error’. Problem: {problem} Solution: {solution}
Appendix D Examples of function similarity
We provide examples of semantic and structural similarity in Table 9. Specifically, functions that share these similarities or have the same docstring are grouped into the same function set.
Semantic Similarity
def gcd(a, b):
␣␣␣␣Findthegreatestcommondivisoroftwonumbers.
␣␣␣␣Args:
␣␣␣␣a(int):Thefirstnumber.
␣␣␣␣b(int):Thesecondnumber.
␣␣␣␣Returns:
␣␣␣␣int:TheGCDofaandb.
while b:
a, b = b, a % b
return a
Structure Similarity
def matrix_transformation(projection_matrix1, projection_matrix2):
␣␣␣␣Calculatetransformationmatrixbymultiplying
␣␣␣␣projectionmatrices.
␣␣␣␣Args:
␣␣␣␣-projection_matrix1:Thefirstprojectionmatrix.
␣␣␣␣-projection_matrix2:Thesecondprojectionmatrix.
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣numpyarray:Thetransformationmatrix.
return np.dot(projection_matrix2, projection_matrix1)
Table 9: Examples of Semantic and Structure Similarity.
Appendix E Format Example for Problem Decomposition